# Heterogeneous Low-Rank Approximation for Federated Fine-tuning of On-Device Foundation Models

### 12 Jan 2024 v1

@@ Abstract
Large foundation models (FMs) adapt surprisingly well to specific domains or tasks with fine-tuning. Federated learning (FL) further enables private FM fine-tuning using the local data on devices. However, the standard FMs’ large size poses challenges for resource-constrained and heterogeneous devices. To address this, we consider FMs with reduced parameter sizes, referred to as on-device FMs (ODFMs). While ODFMs allow on-device inference, computational constraints still hinder efficient federated fine-tuning. We propose a parameter-efficient federated fine-tuning method for ODFMs using heterogeneous low-rank approximations (LoRAs) that addresses system and data heterogeneity. We show that homogeneous LoRA ranks face a trade-off between overfitting and slow convergence, and propose HetLoRA, which employs heterogeneous ranks across clients and eliminates the shortcomings of homogeneous LoRA. By applying rank self-pruning locally and sparsity-weighted aggregation at the server, we combine the advantages of high and low-rank LoRAs, which achieves improved convergence speed and final performance compared to homogeneous LoRA. Furthermore, it offers enhanced computation efficiency compared to full fine-tuning, making it suitable for heterogeneous devices while preserving data privacy.
Heterogeneous Low-Rank Approximation for Federated Fine-tuning of On-Device Foundation Models
@@ 1 Introduction
In our work, we focus on federated fine-tuning, which we define as training a set of parameters collaboratively across clients to obtain a global set of parameters that can be plugged in to the FM for the targeted downstream task. Note that federated fine-tuning is orthogonal to personalization of FMs in federated learning (FL) Guo et al. (2023), which aims to train parameters that perform well for individual clients rather than general downstream tasks. We also define on-device FMs (ODFMs) as models with fewer than several billion parameters that are able to fit into memory on limited capacity clients considering current hardwares.
@@
Federated fine-tuning of ODFMs entails unique challenges that are not present in either standard PEFT of FMs or the standard federated training of models that are not FMs. First, FMs have their zero/few-shot learning capability often supported by their large parameter space that is trained on massive data. However, as shown in Table 1 and previous literature Kojima et al. (2022), FMs’ performance can deteriorate as their sizes get smaller and federated fine-tuning may not merely be useful but even inevitable for ODFMs to perform well for downstream tasks on devices.
Moreover, devices have limited and heterogeneous system capabilities Wang et al. (2019); Bonawitz et al. (2016) and data-heterogeneity Sahu et al. (2020). A suitable PEFT method that flexibly adapts to the heterogeneous devices should be devised for federated fine-tuning of ODFMs. Previous work evaluated PEFT with FL via performing a general evaluation over different PEFT methods naïvely combined with FL Guo et al. (2022); Zhang et al. (2023d); Chen et al. (2022); Wortsman et al. (2023); Yu et al. (2023). However, they do not investigate the practical setting for ODFMs where PEFT methods are catered to heterogeneous system-capabilities and data such as by applying different PEFT strategies to different clients.
In our work. we focus on one of the most prominent PEFT methods, LoRA which proposes to train low-rank approximations of the original model. Using LoRA, the number of trainable parameters is greatly reduced to at most $0.02\%$ of the original ODFM size (see Table 2). The simplest way to apply LoRA to federated fine-tuning is training with homogeneous rank $r$ across the clients as one would train any global model with FL. However, this does not cater to the heterogeneity in FL, where it is even difficult to choose the right LoRA rank for resource limited mobile devices with natural system and data heterogeneity.
@@ 2 Related Work
Federated Fine-Tuning. Recently, interest in the intersection of foundation models and FL has notably increased Zhou et al. (2023); Yu et al. (2023). Many recent work has proposed to simply combine the PEFT methods devised for the centralized setting to FL such as training prompts or adapters collaboratively with FL Guo et al. (2022); Chen et al. (2022); Zhang et al. (2023a); Shysheya et al. (2023); Legate et al. (2023). Another line of work has proposed to perform a few-shot or nearly zero-shot training of FMs with FL for improved communication-efficiency Wortsman et al. (2023); Zhang et al. (2023d). However, these work either overlooks that most devices do not have the resource to fit a general FM Touvron et al. (2023); Brown et al. (2020) (>8B parameters) even for inference or does not consider the heterogeneous system capacities of the clients. It is detrimental to consider these factors since FMs that can actually fit to the devices in FL are usually much smaller making them weaker in the general intelligence capabilities, and also heterogeneous system capacities may prohibit deploying the same sized PEFT parameters across clients.
@@ 3.1 Preliminaries
Formally, we define the pre-trained ODFM as $\mathbf{W}_{0}\in\mathbb{R}^{d\times l}$ and the trainable low-rank decomposed matrix as $\Delta\mathbf{W}\in\mathbb{R}^{d\times l}$. In standard LoRA Hu et al. (2021), the low-rank decomposition of $\Delta\mathbf{W}$ is constructed such that $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$ where $\mathbf{B}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times l}$ are the low rank decomposition of $\Delta\mathbf{W}$ with identical rank $r$ in the centralized setting. Now, let us consider LoRA for federated fine-tuning where there are $M$ total clients. Each client $k\in[M]$ has private data $\mathcal{B}_{k}$ and its corresponding local empirical loss function $F_{k}(\mathbf{W})=\frac{1}{|\mathcal{B}_{k}|}\sum_{\xi\in\mathcal{B}_{k}}\ell(\mathbf{W},\xi)$, where $\ell(\mathbf{W},\xi)$ is the loss for model $\mathbf{W}$ at data sample $\xi$. The optimization task for federated fine-tuning is to collaboratively find the global parameters which we define as $\overline{\mathbf{B}}$ and $\overline{\mathbf{A}}$, given the pretrained knowledge $\mathbf{W}_{0}$ that can minimize the global objective $F(\overline{\mathbf{W}})=\frac{1}{M}\sum_{k=1}^{M}F_{k}(\overline{\mathbf{W}})$ where $\overline{\mathbf{W}}=\mathbf{W}_{0}+\overline{\mathbf{B}}\leavevmode\nobreak\ \overline{\mathbf{A}}$. Throughout the paper, we denote the truncation of a matrix with the :::: symbol for each row and column at the subscript. For instance, if there is no truncation at the row but truncation from $x<l$ at the column for the matrix $\mathbf{W}\in\mathbb{R}^{d\times l}$, we keep all the columns until $x$ and omit the last $l-x$ columns and denote the resulting matrix it as $\mathbf{W}_{:,:x}$.
@@ 3.2 Naïve Case: Homogeneous LoRA
A straightforward way to perform federated fine-tuning with LoRA is to train the LoRA modules $\mathbf{B},\leavevmode\nobreak\ \mathbf{A}$ with homogeneous rank $r$ across all clients with standard FL McMahan et al. (2017). Specifically, first the clients have the pre-trained ODFM weights $\mathbf{W}_{0}$ stored in their devices prior to training for the forward pass when training the LoRA modules. Then, the server sends the global LoRA modules $\overline{\mathbf{B}}^{(t)},\leavevmode\nobreak\ \overline{\mathbf{A}}^{(t)}$ to the set of $m$ selected clients $\mathcal{S}^{(t)}$ per communication round $t$. Each selected client $k\in\mathcal{S}^{(t)}$ trains the LoRA modules on their local data for a few local iterations (usually with mini-batch SGD) and send the updated modules $\mathbf{B}_{k}^{(t)},\leavevmode\nobreak\ \mathbf{A}_{k}^{(t)}$ back to the server. The server then finally updates the global LoRA modules accordingly to $\overline{\mathbf{B}}^{(t+1)}=\sum_{k\in\mathcal{S}^{(t)}}\mathbf{B}_{k}^{(t)}/m,\leavevmode\nobreak\ \overline{\mathbf{A}}^{(t+1)}=\sum_{k\in\mathcal{S}^{(t)}}\mathbf{A}_{k}^{(t)}/m$ and send back to the next set of selected clients for the next communication round. This training process is nearly identical to the standard FL algorithm McMahan et al. (2017) except that the pretrained weights $\mathbf{W}_{0}$ are freezed and locally stored in the clients’ devices and only the LoRA moduels are trained and communicated.
Instead of such homogeneous rank deployment, however, for federated fine-tuning it is not only possible but more practical to consider heterogeneous rank deployment where each client trains different rank LoRA modules. Such setting is motivated and often required from the system constraints of the clients Wang et al. (2021) where most of the clients are only capable of having smaller ranks while a few can handle larger ranks. However, with heterogeneous ranks across for clients, it is non-trivial how to aggregate the heterogenenous LoRA modules at the server and then re-distribute them back to the next set of selected clients. Pushing the limits beyond homogeneous LoRA, we propose HetLoRA in the next section.
Figure 1: Example of (a): the pretrained weights $\mathbf{W}_{0}$ that are stored in the clients’ devices and (b): the heterogeneous ranks assigned to different clients with $r_{\text{min}}=r_{1}<r_{2}<r_{3}=r_{\text{max}}$ for heterogeneous LoRA.
@@ 3.3 Proposed Method: Heterogeneous LoRA
Overview. Recall in heterogeneous LoRA, different clients have different ranks initially assigned to them prior to training to match their computation resources. Our proposed heterogeneous LoRA method, namely HetLoRA, does not assume any specific method to assign the ranks to the clients since we believe that the ranks are not to be decided by the server but rather by the clients and their available resources. We formulate that each client $k$ has a rank $r_{k}$ that is initially assigned to the client where the ranks range from a minimum and maximum rank, i.e., $r_{k}\in[r_{\text{min}},\leavevmode\nobreak\ r_{\text{max}}],\leavevmode\nobreak\ \forall k$ (see Fig. 1). In our proposed HetLoRA there are three major steps for federated fine-tuning of heterogeneous rank LoRA modules: 1) Distribution via Truncation, 2) Local Training with Rank Self-Pruning, and 3) Sparsity-Weighted Aggregation of the LoRA modules. We go into the details of each step in the subsequent paragraphs in this section. An overview of HetLoRA is illustrated in Fig. 2.
@@
1) Distribution via Truncation. First, for each communication round $t$ the server has its initial global LoRA modules $\overline{\mathbf{B}}^{(t)},\leavevmode\nobreak\ \overline{\mathbf{A}}^{(t)}$ with a global rank $r^{(t)}$. The value of the global rank $r^{(t)}$ depends on how we aggregate the heterogeneous rank LoRA modules which we will elaborate on in step 3). The server has to distribute these global LoRA modules to a subset of selected set of clients $\mathcal{S}^{(t)}$ with heterogeneous ranks $r_{k}^{(t)},k\in\mathcal{S}^{(t)}$ for local training [^1] There is a superscript $t$ for the ranks $r_{k}^{(t)}$ across clients which indicates that in HetLoRA these heterogeneous ranks can be changed over the communication rounds via self-pruning explained in step 2.. With the given global LoRA modules, we consider a simple and intuitive method of truncation where the server sends $\overline{\mathbf{B}}^{(t)}_{:,:r_{k}},\leavevmode\nobreak\ \overline{\mathbf{A}}^{(t)}_{:r_{k},:}$ to each client $k$ with rank $r_{k}^{(t)}$ for local training where we omitted the superscript for $r_{k}$ for simplicity.
@@
2) Local Training with Rank Self-Pruning. Initializing their local LoRA modules with the ones received from the server as $\mathbf{B}_{k}^{(t,0)}=\overline{\mathbf{B}}^{(t)}_{:,:r_{k}},\leavevmode\nobreak\ \mathbf{A}_{k}^{(t,0)}=\overline{\mathbf{A}}^{(t)}_{:r_{k},:}$, each client $k\in\mathcal{S}^{(t)}$ can perform $\tau$ local iterations of mini-batch SGD on their local data to minimize the local objective $\frac{1}{|\mathcal{B}_{k}|}\sum_{\xi\in\mathcal{B}_{k}}\ell((\mathbf{B}_{k},\leavevmode\nobreak\ \mathbf{A}_{k}),\xi|\mathbf{W}_{0})$ to send back the updated LoRA modules $\mathbf{B}_{k}^{(t,\tau)}\in\mathbb{R}^{d\times r_{k}^{(t)}}$ and $\mathbf{A}_{k}^{(t,\tau)}\in\mathbb{R}^{r_{k}^{(t)}\times l}$. This is the same process as the standard local training step in vanilla FedAvg McMahan et al. (2017). However, we improve this vanilla local training step by adding a rank self-pruning mechanism where clients self-prune their respective ranks depending on the magnitude of the model parameters.
@@
Specifically, we add a regularization term to the original local objective to get $\min_{\mathbf{B}_{k},\leavevmode\nobreak\ \mathbf{A}_{k}}\frac{1}{|\mathcal{B}_{k}|}\sum_{\xi\in\mathcal{B}_{k}}\ell((\mathbf{B}_{k},\leavevmode\nobreak\ \mathbf{A}_{k}),\xi|\mathbf{W}_{0})+\lambda\left\|\mathbf{B}_{k,:,r_{k}\gamma:r_{k}}\right\|\left\|\mathbf{A}_{k,r_{k}\gamma:r_{k},:}\right\|$ where $\gamma<1$ is a decay-factor that determines how aggressively we want to prune the ranks to a smaller value. The regularization term aims to minimize the norm of the last few ranks, which will become smaller if the first loss term $\frac{1}{|\mathcal{B}_{k}|}\sum_{\xi\in\mathcal{B}_{k}}\ell((\mathbf{B}_{k},\leavevmode\nobreak\ \mathbf{A}_{k}),\xi|\mathbf{W}_{0})$ is not very large. After training with the new local objective we compare the norm of the updated LoRA modules’ last few layers $\left\|\mathbf{B}_{k,:,r_{k}\gamma:r_{k}}\right\|\left\|\mathbf{A}_{k,r_{k}\gamma:r_{k},:}\right\|$ with the ones from the initially received LoRA modules. If the former is smaller we prune the last few layers (which pruning intensity is determined by $\gamma$) and send back the LoRA modules with a smaller rank. This means that for the LoRA modules which incurs a small local loss, i.e., well-trained on the clients’ local data, the LoRA modules are more likely to be pruned to a smaller rank. Such pruning allows HetLoRA to reduce the noise in the LoRA modules introduced by clients having a larger rank than the actual rank that their data complexity requires and also reduce the model complexity of the LoRA modules to improve generalization and prevent overfitting (see Table 4). Once the rank is pruned for a client, the client saves the updated rank and uses it as the starting rank if selected for future communication rounds. The client then sends back their updated and possibly rank-pruned LoRA modules to the server for the modules to be processed in the next step, server-side aggregation.
@@
3) Sparsity-Weighted Aggregation. Finally, the last step of HetLoRA is aggregating the received heterogeneous LoRA modules $\mathbf{B}_{k}^{(t,\tau)},\leavevmode\nobreak\ \mathbf{A}_{k}^{(t,\tau)},k\in\mathcal{S}^{(t)}$. A straightforward way to aggregate the hetergeneous LoRA modules is using zero-padding to all the received LoRA modules with $r_{i}^{(t)}<\max\{r_{k}^{(t)}|k\in\mathcal{S}^{(t)}\}$ and then perform simple averaging over the modules. However, such naive aggregation can lead to biasing the model towards higher rank clients even when these clients may not hold valuable training information, i.e., have low data complexity, giving noisy updates.
@@
Based on this insight we propose a sparsity-weighted aggregation scheme where the server reconstructs these LoRA modules to the full model as $\Delta\mathbf{W}_{k}^{(t)}=\mathbf{B}_{k}^{(t)}\mathbf{A}_{k}^{(t)}$ and gets the norm of the singular value vectors from the full models denoted as $\mathbf{S}_{k}^{(t)}$ by calculating $\|\Delta\mathbf{W}_{k}^{(t)}\|_{F}$. Note that the costly process of performing SVD for each of the full model $\Delta\mathbf{W}_{k}^{(t)}$ can be avoided by simply calculating the Frobenius norm of $\Delta\mathbf{W}_{k}^{(t)}$ (see Lemma 1.2 in Guruswami and Kannan (2012)). The server then weighs the LoRA modules accordingly by the norm of the singular value vectors, updating the global LoRA modules as $\overline{\mathbf{B}}^{(t+1)}=\sum_{k\in\mathcal{S}^{(t)}}\|\mathbf{S}_{k}^{(t)}\|\mathbf{B}_{k}^{(t)}/\mathbf{Z}^{(t)},\leavevmode\nobreak\ \overline{\mathbf{A}}^{(t+1)}=\sum_{k\in\mathcal{S}^{(t)}}\|\mathbf{S}_{k}^{(t)}\|\mathbf{A}_{k}^{(t)}/\mathbf{Z}^{(t)}$ where $\mathbf{Z}^{(t)}\vcentcolon=\sum_{k\in\mathcal{S}^{(t)}}\|\mathbf{S}_{k}^{(t)}\|$ is the normalizing factor. This way, we can de-emphasize the larger rank assigned-clients that have rather less informative updates, and more emphasize the smaller rank assigned-clients that have more informative ones.
3.4 Why not Simply Reconstruct First,
then Redistribute the LoRA modules?

### 20 Feb 2024 v2

# Heterogeneous LoRA for Federated Fine-tuning of On-Device Foundation Models

@@ Abstract
Foundation models (FMs) adapt well to specific domains or tasks with fine-tuning, and federated learning (FL) enables the potential for privacy-preserving fine-tuning of the FMs with on-device local data. For federated fine-tuning of FMs, we consider the FMs with small to medium parameter sizes of single digit billion at maximum, referred to as on-device FMs (ODFMs) that can be deployed on devices for inference but can only be fine-tuned with parameter efficient methods. In our work, we tackle the data and system heterogeneity problem of federated fine-tuning of ODFMs by proposing a novel method using heterogeneous low-rank approximations (LoRAs), namely HetLoRA. First, we show that the naive approach of using homogeneous LoRA ranks across devices face a trade-off between overfitting and slow convergence, and thus propose HetLoRA, which allows heterogeneous ranks across client devices and efficiently aggregates and distributes these heterogeneous LoRA modules. By applying rank self-pruning locally and sparsity-weighted aggregation at the server, HetLoRA combines the advantages of high and low-rank LoRAs, which achieves improved convergence speed and final performance compared to homogeneous LoRA. Furthermore, HetLoRA offers enhanced computation efficiency compared to full fine-tuning, making it suitable for federated fine-tuning across heterogeneous devices.
Heterogeneous LoRA for Federated Fine-tuning of
On-Device Foundation Models
@@ 1 Introduction
In our work, we focus on such federated fine-tuning scenarios, where we train a set of parameters collaboratively across clients to obtain a global set of parameters that can be plugged in to the FM for the targeted downstream task. Note that federated fine-tuning is orthogonal to personalization of FMs in federated learning (FL) Guo et al. (2023), which aims to train parameters that perform well for individual clients rather than general downstream tasks. We also define on-device FMs (ODFMs) as models with few billion parameters at max that are able to fit into memory on limited capacity clients considering current hardwares.
@@
Federated fine-tuning of ODFMs entails unique challenges non-present in either the standard PEFT of FMs or the standard federated training of models that are not FMs. First, FMs have their zero/few-shot learning capability often supported by their large parameter space that is trained on massive data. However, as we show in Table 1 and also presented by previous literature Kojima et al. (2022), FMs’ performance deteriorates as their sizes get smaller and federated fine-tuning may not merely be useful but inevitable for ODFMs to perform well for downstream tasks on devices.
Moreover, devices have limited and heterogeneous system capabilities Wang et al. (2019); Bonawitz et al. (2016) and data distributions Sahu et al. (2020). A suitable PEFT method that flexibly adapts to such heterogeneity across devices should be investigated for federated fine-tuning of ODFMs. Previous work evaluated PEFT with FL via performing a general evaluation over different PEFT methods naïvely combined with FL Guo et al. (2022); Zhang et al. (2023d); Chen et al. (2022); Wortsman et al. (2023); Yu et al. (2023). However, they do not consider the practical setting for ODFMs where PEFT methods are catered to the system and data heterogeneity of clients.
In our work. we focus on one of the most prominent PEFT methods, LoRA Hu et al. (2021) which proposes to train low-rank approximations of the original model. Using LoRA, the number of trainable parameters is greatly reduced to at most $0.02\%$ of the original ODFM size (see Table 2). The simplest way to apply LoRA to federated fine-tuning is training with homogeneous rank $r$ across the clients as one would train any global model with FL. However, this does not cater to the heterogeneity in FL, where it is even difficult to choose the right LoRA rank for resource limited mobile devices with natural system and data heterogeneity.
Figure 1: Overview of heterogeneous rank deployment of LoRA: the pretrained weights $\mathbf{W}_{0}$ are stored on-device and heterogeneous ranks are assigned to different clients with $r_{\text{min}}=r_{1}<r_{2}<r_{3}=r_{\text{max}}$. In our proposed HetLoRA, the server receives the trained heterogeneous LoRA modules and aggregates them with sparsity-weighted aggregation to update the global LoRA module.
@@ 2 Related Work
Federated Fine-Tuning. Recently, interest in the intersection of FMs and FL has notably increased Zhou et al. (2023); Yu et al. (2023). Many recent work has proposed to combine the PEFT methods devised for the centralized setting to FL such as training prompts or adapters collaboratively with FL Guo et al. (2022); Chen et al. (2022); Zhang et al. (2023a); Shysheya et al. (2023); Legate et al. (2023). Another line of work has proposed to perform a few-shot or nearly zero-shot training of FMs with FL for improved communication-efficiency Wortsman et al. (2023); Zhang et al. (2023d). However, these work either overlooks that most devices do not have the resource to fit a general FM Touvron et al. (2023); Brown et al. (2020) (>8B parameters) even for inference or does not consider the heterogeneous system capacities of the clients. It is detrimental to consider these factors since FMs that actually fits to the devices in FL are much smaller, making them weaker in the general intelligence capabilities, and also heterogeneous system capacities may prohibit deploying same sized PEFT parameters across clients.
@@ 3.1 Preliminaries
Formally, we define the pre-trained ODFM as $\mathbf{W}_{0}\in\mathbb{R}^{d\times l}$ and the trainable low-rank decomposed matrix as $\Delta\mathbf{W}\in\mathbb{R}^{d\times l}$. In standard LoRA Hu et al. (2021) under the centralized setting, the low-rank decomposition of $\Delta\mathbf{W}$ is constructed such that $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$ where $\mathbf{B}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times l}$ are the low rank decomposition of $\Delta\mathbf{W}$ with identical rank $r$. Now, let us consider LoRA for federated fine-tuning where there are $M$ total clients. Each client $k\in[M]$ has private data $\mathcal{B}_{k}$ and its corresponding local empirical loss function $F_{k}(\mathbf{W})=\frac{1}{|\mathcal{B}_{k}|}\sum_{\xi\in\mathcal{B}_{k}}\ell(\mathbf{W},\xi)$, where $\ell(\mathbf{W},\xi)$ is the loss for model $\mathbf{W}$ at data sample $\xi$. The optimization task for federated fine-tuning is to collaboratively find the global parameters which we define as $\overline{\mathbf{B}}$ and $\overline{\mathbf{A}}$, given the pretrained knowledge $\mathbf{W}_{0}$ that can minimize the global objective $F(\overline{\mathbf{W}})=\frac{1}{M}\sum_{k=1}^{M}F_{k}(\overline{\mathbf{W}})$ where $\overline{\mathbf{W}}=\mathbf{W}_{0}+\overline{\mathbf{B}}\leavevmode\nobreak\ \overline{\mathbf{A}}$. Later in the paper, when introducing heterogeneous LoRA we truncate the LoRA modules’ rank dimension, for example from $\mathbf{B}\in\mathbb{R}^{d\times r},\leavevmode\nobreak\ \mathbf{A}\in\mathbb{R}^{r\times l}$ to $\mathbf{B}^{\prime}\in\mathbb{R}^{d\times r^{\prime}},\leavevmode\nobreak\ \mathbf{A}^{\prime}\in\mathbb{R}^{r^{\prime}\times l}$ where $r^{\prime}<r$. Throughout the paper, we denote such truncation of a matrix with the :::: symbol for each row and column at the subscript. For instance, for truncation to $r^{\prime}<r$ at the column for the matrix $\mathbf{B}\in\mathbb{R}^{d\times r}$, we keep all the columns until $r^{\prime}$ and omit the last $r-r^{\prime}$ columns and denote the resulting matrix it as $\mathbf{B}_{:,:r^{\prime}}$.
@@ 3.2 Naïve Case: Homogeneous LoRA
A straightforward way to perform federated fine-tuning with LoRA is to train the LoRA modules $\mathbf{B},\leavevmode\nobreak\ \mathbf{A}$ with homogeneous rank $r$ across all clients with standard FL McMahan et al. (2017). Specifically, first the clients have the pre-trained ODFM weights $\mathbf{W}_{0}$ stored in their devices prior to training for the forward pass when training the LoRA modules. Then, the server sends the global LoRA modules $\overline{\mathbf{B}}^{(t)},\leavevmode\nobreak\ \overline{\mathbf{A}}^{(t)}$ to the set of $m$ selected clients $\mathcal{S}^{(t)}$ per communication round $t$. Each selected client $k\in\mathcal{S}^{(t)}$ trains the LoRA modules on their local data for a few local iterations (usually with mini-batch SGD) and send the updated modules $\mathbf{B}_{k}^{(t)},\leavevmode\nobreak\ \mathbf{A}_{k}^{(t)}$ back to the server. The server then updates the global LoRA modules accordingly to $\overline{\mathbf{B}}^{(t+1)}=\sum_{k\in\mathcal{S}^{(t)}}\mathbf{B}_{k}^{(t)}/m,\leavevmode\nobreak\ \overline{\mathbf{A}}^{(t+1)}=\sum_{k\in\mathcal{S}^{(t)}}\mathbf{A}_{k}^{(t)}/m$ and sends back to the next set of selected clients for the next communication round. This training process is nearly identical to the standard FL algorithm McMahan et al. (2017) except that the pretrained weights $\mathbf{W}_{0}$ are freezed and locally stored in the clients’ devices and only the LoRA moduels are trained and communicated.
Instead of such homogeneous rank deployment across all clients, it is not only possible but more practical to use heterogeneous rank deployment for federated fine-tuning. This involves training LoRA modules with varying ranks on different clients, based on their system capabilities. Such setting is motivated and often required from the system constraints of the clients Wang et al. (2021) where most of the clients are only capable of having smaller ranks while a few can handle larger ranks. However, this approach poses challenges in aggregating and redistributing the LoRA modules. To address these challenges, we introduce a solution called HetLoRA, which pushes the limits beyond homogeneous LoRA deployment.
@@ 3.3 Proposed Method: Heterogeneous LoRA
Overview. Our proposed heterogeneous LoRA method, namely HetLoRA, is not restricted to any specific method to assign the ranks to the clients and the clients can decide their respective ranks themselves. For formality, in our paper, we formulate that each client has a rank denoted as $r_{k}$, within a range of $r_{k}\in[r_{\text{min}},\leavevmode\nobreak\ r_{\text{max}}],\leavevmode\nobreak\ \forall k$ (see Fig. 1). HetLoRA comprises three steps: 1) Distribution via Truncation, 2) Local Training with Rank Self-Pruning, and 3) Sparsity-Weighted Aggregation of the LoRA modules. These steps are detailed further in the subsequent paragraphs. An overview of HetLoRA is illustrated in Fig. 2.
@@
1) Distribution via Truncation. At the begining of each communication round $t$, the server holds initial global LoRA modules $\overline{\mathbf{B}}^{(t)},\leavevmode\nobreak\ \overline{\mathbf{A}}^{(t)}$ with a global rank $r^{(t)}$. The value of the global rank $r^{(t)}$ depends on how we aggregate the heterogeneous rank LoRA modules which is elaborated on in step 3). The server then distributes these global LoRA modules to a subset of selected set of clients $\mathcal{S}^{(t)}$ with heterogeneous ranks $r_{k}^{(t)},k\in\mathcal{S}^{(t)}$ for local training [^1] There is a superscript $t$ for the ranks $r_{k}^{(t)}$ across clients which indicates that in HetLoRA these heterogeneous ranks can be changed over the communication rounds via self-pruning explained in step 2).. With the given global LoRA modules, we consider a simple and intuitive method of truncation where the server sends $\overline{\mathbf{B}}^{(t)}_{:,:r_{k}},\leavevmode\nobreak\ \overline{\mathbf{A}}^{(t)}_{:r_{k},:}$ to each client $k$ with rank $r_{k}^{(t)}$ for local training where we omitted the superscript for $r_{k}$ for simplicity.
@@
2) Local Training with Rank Self-Pruning. After receiving LoRA modules from the server as $\mathbf{B}_{k}^{(t,0)}=\overline{\mathbf{B}}^{(t)}_{:,:r_{k}},\leavevmode\nobreak\ \mathbf{A}_{k}^{(t,0)}=\overline{\mathbf{A}}^{(t)}_{:r_{k},:}$, each client $k\in\mathcal{S}^{(t)}$ performs $\tau$ local iterations of mini-batch SGD on their local data to minimize the local objective $\frac{1}{|\mathcal{B}_{k}|}\sum_{\xi\in\mathcal{B}_{k}}\ell((\mathbf{B}_{k},\leavevmode\nobreak\ \mathbf{A}_{k}),\xi|\mathbf{W}_{0})$, and sends back the updated LoRA modules $\mathbf{B}_{k}^{(t,\tau)}\in\mathbb{R}^{d\times r_{k}^{(t)}}$ and $\mathbf{A}_{k}^{(t,\tau)}\in\mathbb{R}^{r_{k}^{(t)}\times l}$ to the server. This is the same process as the standard local training step in vanilla FedAvg McMahan et al. (2017). However, we improve this vanilla local training step by adding a rank self-pruning mechanism where clients self-prune their respective ranks depending on the magnitude of the model parameters.
@@
Specifically, we add a regularization term to the original local objective to get $\min_{\mathbf{B}_{k},\leavevmode\nobreak\ \mathbf{A}_{k}}\frac{1}{|\mathcal{B}_{k}|}\sum_{\xi\in\mathcal{B}_{k}}\ell((\mathbf{B}_{k},\leavevmode\nobreak\ \mathbf{A}_{k}),\xi|\mathbf{W}_{0})+\lambda\left\|\mathbf{B}_{k,:,r_{k}\gamma:r_{k}}\right\|\left\|\mathbf{A}_{k,r_{k}\gamma:r_{k},:}\right\|$ where $\gamma<1$ is a decay-factor that determines how aggressively we want to prune the ranks to a smaller value. The regularization term aims to minimize the norm of the last few ranks, which will become smaller if the first loss term $\frac{1}{|\mathcal{B}_{k}|}\sum_{\xi\in\mathcal{B}_{k}}\ell((\mathbf{B}_{k},\leavevmode\nobreak\ \mathbf{A}_{k}),\xi|\mathbf{W}_{0})$ is not very large. After training with the new local objective we compare the norm of the updated LoRA modules’ last few layers $\left\|\mathbf{B}_{k,:,r_{k}\gamma:r_{k}}\right\|\left\|\mathbf{A}_{k,r_{k}\gamma:r_{k},:}\right\|$ with the ones from the initially received LoRA modules. If the former is smaller we prune the last few layers (pruning intensity is determined by $\gamma$) and send back the LoRA modules with a smaller rank. This means that for the LoRA modules which incurs a small local loss, i.e., well-trained on the clients’ local data, the LoRA modules are more likely to be pruned to a smaller rank.
Such pruning allows HetLoRA to reduce the noise in the LoRA modules introduced by clients having a larger rank than the actual rank that their data complexity requires, and also reduces the complexity of the LoRA modules to improve generalization and prevent overfitting (see Table 4). Once the rank is pruned for a client, the client saves the updated rank and uses it as the starting rank if selected for future communication rounds. The client then sends back their updated and possibly rank-pruned LoRA modules to the server for the modules to be processed in the next server-side aggregation step.
@@
3) Sparsity-Weighted Aggregation. Finally, the last step of HetLoRA is aggregating the received heterogeneous LoRA modules $\mathbf{B}_{k}^{(t,\tau)},\leavevmode\nobreak\ \mathbf{A}_{k}^{(t,\tau)},k\in\mathcal{S}^{(t)}$. A straightforward way to aggregate the hetergeneous LoRA modules is using zero-padding to all the received LoRA modules with $r_{i}^{(t)}<\max\{r_{k}^{(t)}|k\in\mathcal{S}^{(t)}\}$ and then perform simple averaging over the modules. However, such naive aggregation can lead to biasing the model towards higher rank clients even when these clients may not hold valuable training information, i.e., having low data complexity, giving noisy updates.
@@
Based on this insight we propose a sparsity-weighted aggregation scheme where the server reconstructs these LoRA modules to the full model as $\Delta\mathbf{W}_{k}^{(t)}=\mathbf{B}_{k}^{(t)}\mathbf{A}_{k}^{(t)}$ and gets the norm of the singular value vectors from the full models denoted as $\mathbf{S}_{k}^{(t)}$ by calculating $\|\Delta\mathbf{W}_{k}^{(t)}\|_{F}$. Note that the costly process of performing SVD for each of the full model $\Delta\mathbf{W}_{k}^{(t)}$ can be avoided by simply calculating the Frobenius norm of $\Delta\mathbf{W}_{k}^{(t)}$ (see Lemma 1.2 in Guruswami and Kannan (2012)). The server then weighs the LoRA modules with aggregation weight $p_{k}^{(t)}$ which is proportional to the norm of the singular value vectors. Formally, we have the the global LoRA modules updated as $\overline{\mathbf{B}}^{(t+1)}=\sum_{k\in\mathcal{S}^{(t)}}p_{k}^{(t)}\mathbf{B}_{k}^{(t)},\leavevmode\nobreak\ \overline{\mathbf{A}}^{(t+1)}=\sum_{k\in\mathcal{S}^{(t)}}p_{k}^{(t)}\mathbf{A}_{k}^{(t)}$ where $p_{k}^{(t)}\vcentcolon=\|\mathbf{S}_{k}^{(t)}\|/\mathbf{Z}^{(t)}$ with normalizing factor $\mathbf{Z}^{(t)}\vcentcolon=\sum_{k^{\prime}\in\mathcal{S}^{(t)}}\|\mathbf{S}_{k^{\prime}}^{(t)}\|$. This way, we can de-emphasize the larger rank assigned-clients that have rather less informative updates, and more emphasize the smaller rank assigned-clients that have more informative ones.
3.4 Why not Simply Reconstruct First, then Redistribute the LoRA modules?